# Failure of famous functional inequalities on Finsler manifolds: the influence of $S$-curvature

### v1

@@ Abstract.
The validity of functional inequalities on Finsler metric measure manifolds is based on three non-Riemannian quantities, namely, the reversibility, flag curvature and $S$-curvature induced by the measure. Under mild assumptions on the reversibility and flag curvature, it turned out that famous functional inequalities – as Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality – usually hold on forward complete Finsler manifolds with non-positive $S$-curvature, cf. Huang, Kristály and Zhao [Trans. Amer. Math. Soc., 2020]. In this paper however we prove that – under similar assumptions on the reversibility and flag curvature as before – the aforementioned functional inequalities fail whenever the $S$-curvature is positive. Accordingly, our results clearly reveal the deep dependence of functional inequalities on the $S$-curvature. As a consequence of these results, we establish surprising analytic aspects of Finsler manifolds: if the flag curvature is non-positive, the Ricci curvature is bounded from below and the $S$-curvature is positive, then the reversibility turns out to be infinite. Examples are presented on general Funk metric spaces, where the $S$-curvature plays again a decisive role.
@@
A. Kristály is supported by the Excellence Researcher Program ÓE-KP-2-2022 of Óbuda University, Hungary. B. Li is supported by Natural Science Foundation of Zhejiang Province (No. LY23A010012), Natural Science Foundation of Ningbo (2021J110) and K.C. Wong Magna Fund in Ningbo University. W. Zhao is supported by Natural Science Foundation of Shanghai (No. 21ZR1418300) and Science and Technology Project of Xinjiang Production and Construction Corps (No. 2023CB008-12).
@@ 1. Introduction
As the title of S.S. Chern’s famous 1996 paper [7] shows, his credo was that ‘Finsler geometry is just Riemannian geometry without the quadratic restriction’. In addition, Chern concluded his paper by claiming that ”(…) almost all the results of Riemannian geometry can be developed in the Finsler setting. It is remarkable that one needs only some conceptual adjustment, no essential new ideas being necessary.” To a certain point, Chern’s position cannot be discredited, since most of the well-known results from Riemannian geometry – by suitable adaptations – have their natural Finslerian counterparts, e.g., Hopf–Rinow, Cartan–Hadamard, Gauss–Bonnet and Bonnet–Myers theorems as well as Rauch, Bishop–Gromov and Bishop–Günther comparison principles; see Auslander [1], Bao, Chern and Shen [2], Rund [25] and Shen [27]. However, the development of Finsler geometry over the last three decades clearly shows that fundamentally new ideas are needed to solve deep problems, revealing in this way genuine differences between Riemannian and Finsler geometries. In support of this thesis, we recall some open questions and spectacular results from Finsler geometry, the solutions to which are far from being tied to the Riemannian world:
  • Classification of Finsler manifolds with constant flag curvature. Unlike the Hopf classification in Riemannian geometry, no full characterization is available for Finsler manifolds having constant flag curvature; in fact, it seems that different subclasses of Finsler manifolds – as Minkowski, Berwald, Landsberg, projectively flat and Randers spaces – play a crucial role in such a characterization, see e.g., Bao, Robles and Shen [3], Li [19], and Shen [28, 29, 31].
  • Busemann NPC spaces. Busemann non-positively curved (shortly, NPC) spaces on Riemannian manifolds are characterized by the non-positivity of the sectional curvature, see Busemann [4] and Jost [13]. It was shown however by Kelly and Straus [15] that any strictly convex domain $D\subset\mathbb{R}^{2}$ endowed with the Hilbert metric (which is a projectively flat Finsler manifold with constant negative curvature) is a Busemann NPC space if and only if the boundary of $D$ is an ellipse and the Hilbert geometry is hyperbolic. Ivanov and Lytchak [12] proved that a Finsler manifold is a Busemann NPC space if and only if it is a Berwald manifold of non-positive flag curvature.
  • Sobolev spaces over Finsler manifolds. Contrary to the Riemannian setting, the Sobolev spaces over noncompact, irreversible Finsler manifolds should not have even a vector space structure, see Kristály and Rudas [18]. This construction was carried out over the Funk metric defined on the $n$-dimensional Euclidean unit ball $\mathbb{B}^{n}$.
@@
Let $(M,F,\operatorname{\mathfrak{m}})$ be a Finsler metric measure manifold ($\operatorname{\mathsf{FMMM}}$, for short), i.e., a Finsler manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$, see Section 2 for details. Huang, Kristály and Zhao [10] already pointed out that three non-Riemannian quantities strongly influence the validity of functional inequalities on $(M,F,\operatorname{\mathfrak{m}})$, namely,
@@
  • $S$-curvature induced by the measure $\operatorname{\mathfrak{m}}$;
  • flag curvature.
@@
It is easy to see that $\lambda_{F}(M)\geq 1$ with equality if and only if $F$ is reversible (i.e., symmetric). Obviously, all Riemannian metrics are reversible, but there are infinitely many irreversible Finsler metrics. Typical examples with infinite reversibility are Funk metric spaces, which served to prove that Sobolev spaces over such Finsler manifolds are not linear structures, see Farkas, Kristály and Varga [9] and Kristály and Rudas [18]. Even if the Finsler metric is explicitly given, it is usually a hard task to determine whether the reversibility is finite or not.
Unlike in the Riemannian case, there is no a canonical measure on Finsler manifolds. Thus, various measures can be introduced whose behavior may be genuinely different. Every smooth positive measure $\operatorname{\mathfrak{m}}$ on a Finsler manifold $(M,F)$ induces two non-Riemannian quantities, which are the so-called distortion $\tau$ and $S$-curvature $\mathbf{S}$, respectively, see Shen [27]. More precisely, set $\operatorname{d\mathfrak{m}}:=\sigma(x){\operatorname{d}}x^{1}\wedge\cdots\wedge{\operatorname{d}}x^{n}$ in some local coordinate system $(x^{i})$. Thus, for any $y\in T_{x}M\backslash\{0\}$, define
@@
where $R_{y}$ is the Riemannian curvature of $F.$ Clearly, the flag curvature naturally defines the Ricci curvature ${\bf Ric}.$ As expected, on any Riemannian manifold endowed with its canonical measure the $S$-curvature vanishes, while the flag curvature reduces to the sectional curvature.
Before stating our main results, we recall the main achievements from Huang, Kristály and Zhao [10], which represent the starting points of our present study; in fact, they roughly state that if
@@
    $\ \ \ \ \ {\displaystyle\left(\int_{M}\max\left\{{F}^{*2}({\pm\rm d}u)\right\}\operatorname{d\mathfrak{m}}\right)}\left(\int_{M}{r^{2}u^{2}}\operatorname{d\mathfrak{m}}\right)\geq\frac{n^{2}}{4}{\displaystyle\left(\int_{M}{u^{2}}\operatorname{d\mathfrak{m}}\right)^{2}},\ \forall u\in C^{\infty}_{0}(M);$ (1.3)
@@
  • Caffarelli–Kohn–Nirenberg interpolation inequality; see [10, Theorem 1.1]. If $0<q<2<p$ and $2<n<2(p-q)/(p-2)$, then for every $u\in C^{\infty}_{0}(M)$ one has
@@
On the other hand, if $(M,F,\operatorname{\mathfrak{m}})$ is an $n$-dimensional reversible $\operatorname{\mathsf{FMMM}}$, and the ’complementary’ case to (1.1) holds, i.e.,
@@
The above set of results contain in certain sense the genesis of our first main result – which reads as follows – and roughly shows that the validity/failure of functional inequalities on Finsler manifolds is decided by the sign of the $S$-curvature:
@@
where $k,h$ are two constants with $h>k\geq 0$. Then the following statements hold::::
@@
    $\inf_{u\in C^{\infty}_{0}(M)}\frac{\displaystyle\int_{M}{F}^{*p}({\rm d}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}{|u|^{p}}\operatorname{d\mathfrak{m}}}=0;$
@@
    $\inf_{u\in C^{\infty}_{0}(M)}\frac{\displaystyle\int_{M}{F}^{*p}({\rm d}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}=0;$
@@
    $\inf_{u\in C^{\infty}_{0}({M})\backslash\{0\}}\frac{\left(\displaystyle\int_{{M}}F^{*2}({\rm d}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{{M}}r^{2}u^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{{M}}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}=0;$
@@
    the Caffarelli–Kohn–Nirenberg interpolation inequality fails on $(M,F,\operatorname{\mathfrak{m}})$, i.e., if $0<q<2<p$ and $2<n<2(p-q)/(p-2)$, then
    $\inf_{u\in C^{\infty}_{0}({M})\backslash\{0\}}\frac{\left(\displaystyle\int_{{M}}F^{*2}({\rm d}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{{M}}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{{M}}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}=0.$
Intuitively, the $S$-curvature describes the change of measure along geodesics, see Shen [30]. As far as we know, only a few results are available in the literature concerning the impact of $S$-curvature on global properties of Finsler manifolds. In the light of the results from [10], Theorem 1.1 provides a concluding answer about the validity of functional inequalities in Finsler manifolds, which is captured by the behavior of the $S$-curvature. In particular, all the statements except (i) in Theorem 1.1 will be invalid if the condition $h>0$ is relaxed to $h\geq 0$, showing in this way the optimality of our results. Indeed, every reversible Minkowski space endowed with the Busemann–Hausdorff measure is a Cartan–Hadamard manifold with $\mathbf{K}=0$ and $\mathbf{S}=0$ (i.e., $h=k=0$); thus the (sharp) inequalities from (1.2)-(1.4) all are valid, which contradict the conclusions from Theorem 1.1. We also notice that there are large classes of non-Riemannian reversible Minkowski spaces; see Example 6.1.
The proof of Theorem 1.1 is obtained by considering polar coordinates on $M$ with a suitable disintegration of the measure $\operatorname{\mathfrak{m}}$, combined with the key assumption (1.5) with $h>k\geq 0$, comparison principles and suitably chosen test functions (with appropriate approximation arguments). For the proof of (i)-(iii) the Guassian bubble $u_{\alpha}=-e^{-\alpha r}$ is used, while for (iv) a Talentian bubble is considered. The Gaussian bubble has been used in Kristály [17] and Huang, Kristály and Zhao [10] to prove (i) and (ii), respectively, in the special case of the Funk metric on the Euclidean unit ball $\mathbb{B}^{n}$. In general, according to Li [19] and Shen [31, 26], for every non-empty bounded strongly convex domain $\Omega\subset\mathbb{R}^{n}$ (containing the origin $\mathbf{0}$) with smooth boundary, there always exists a Finsler metric $F$, which is a Funk metric, and $(\Omega,F)$ is a Cartan–Hadamard manifold with $\mathbf{K}=-1/4$, see Section 6. Such a pair $(\Omega,F)$ is called a Funk metric space, and it is the model space where our Theorem 1.1 can be applied. In the particular case when $\Omega=\mathbb{B}^{n}$ and $\operatorname{\mathfrak{m}}=\operatorname{\mathfrak{m}}_{BH}$, the Funk metric on $\mathbb{B}^{n}$ is
@@
see Shen [31, 26, 27], for which ${\bf Ric}=-(n-1)/4$ and ${\bf S}_{BH}=(n+1)/2,$ thus $h=\frac{n+1}{2(n-1)}>\frac{1}{2}=k>0.$ However, the symmetrization of $F$ is $\tilde{F}(x,y)=(F(x,y)+F(x,-y))/2,$ which is precisely the Riemannian Hilbert/Klein metric on $\mathbb{B}^{n}$ with ${\bf K}=-1$ (so ${\bf Ric}=-(n-1)$) and ${\bf S}=0$; thus the inequalities (1.2)-(1.4) apply instead of Theorem 1.1.
@@
Obviously, every Funk metric space endowed with the Busemann-Hausdorff measure verifies the assumptions of Theorem 1.2. The additional condition $\mathbf{K}\leq 0$ is mild. In fact, if the flag curvature is bounded from below by a positive number, then the Bonnet–Myers theorem implies the compactness of the manifold, see [2, Theorem 7.7.1], in which case the reversibility is finite.
We recall that another important notion in Finsler geometry is the weighted Ricci curvature $\mathbf{Ric}_{N}$ for $N\in[n,\infty]$. Although $\mathbf{Ric}_{N}$ is a combination of $\mathbf{Ric}$ and $\mathbf{S}$, we show that one cannot replace the separated conditions from (1.5) (or (1.7)) by $\mathbf{Ric}_{N}\geq-K^{2}$ for any $K\geq 0$. In fact, every reversible Minkowski space equipped with the Busemann–Hausdorff measure satisfies $\mathbf{Ric}_{N}=0$, but neither Theorem 1.1 nor Theorem 1.2 holds on such spaces; for further examples, see subsection 6.2. Moreover, it is not appropriate either to assume $\mathbf{Ric}_{N}\geq K>0,$ since in this case the manifold is compact and consequently, its reversibility is finite.
The paper is organized as follows. Section 2 is devoted to preliminaries on Finsler geometry together with some results about volume comparison. In Section 3, we provide some immediate properties of Sobolev spaces over irreversible Finsler manifolds. Here, we also establish an approximation result which will be useful in our main results. The proof of our main Theorem 1.1 is given in Section 4, while the proof of Theorem 1.2 is presented in Section 5. Section 6 is devoted to the study of Funk metric spaces, by providing examples and counterexamples supporting the sharpness of our main results.

### v2

@@ Abstract.
The validity of functional inequalities on Finsler metric measure manifolds is based on three non-Riemannian quantities, namely, the reversibility, flag curvature and $S$-curvature induced by the measure. Under mild assumptions on the reversibility and flag curvature, it turned out that famous functional inequalities – as Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality – usually hold on forward complete Finsler manifolds with non-positive $S$-curvature, cf. Huang, Kristály and Zhao [Trans. Amer. Math. Soc., 2020]. In this paper however we prove that – under similar assumptions on the reversibility and flag curvature as before – the aforementioned functional inequalities fail whenever the $S$-curvature is positive. Accordingly, our results clearly reveal the deep dependence of functional inequalities on the $S$-curvature. As a consequence of these results, we establish analytic aspects of Finsler manifolds, e.g., if the flag curvature is non-positive, the Ricci curvature is bounded from below and the $S$-curvature is positive, then the reversibility turns out to be infinite. Further topological properties and examples are presented on general Funk metric spaces, where the $S$-curvature plays again a decisive role.
@@
A. Kristály is supported by the Excellence Researcher Program ÓE-KP-2-2022 of Óbuda University, Hungary. B. Li is supported by Natural Science Foundation of Zhejiang Province (No. LY23A010012) and Natural Science Foundation of Ningbo (No. 2024J017). W. Zhao is supported by Natural Science Foundation of China (No. 12471045) and Science and Technology Project of Xinjiang Production and Construction Corps (No. 2023CB008-12).
@@ 1. Introduction
As the title of S. S. Chern’s famous 1996 paper [7] shows, his credo was that ‘Finsler geometry is just Riemannian geometry without the quadratic restriction’. In addition, Chern concluded his paper by claiming that “(…) almost all the results of Riemannian geometry can be developed in the Finsler setting. It is remarkable that one needs only some conceptual adjustment, no essential new ideas being necessary.” To some extent, Chern’s position cannot be discredited, since most of the well-known results from Riemannian geometry – by suitable adaptations – have their natural Finslerian counterparts, e.g., Hopf–Rinow, Cartan–Hadamard, Gauss–Bonnet and Bonnet–Myers theorems as well as Rauch, Bishop–Gromov and Bishop–Günther comparison principles; see Auslander [1], Bao–Chern–Shen [2], Rund [26], Shen [28] and Ohta [21]. However, the development of Finsler geometry over the last three decades clearly shows that fundamentally new ideas are needed to solve deep problems, revealing in this way genuine differences between Riemannian and Finsler geometries. In support of this thesis, we recall some open questions and spectacular results from Finsler geometry, the solutions to which are far from being tied to the Riemannian world:
  • Classification of Finsler manifolds with constant flag curvature. Unlike the Hopf classification in Riemannian geometry, no full characterization is available for Finsler manifolds having constant flag curvature; in fact, it seems that different subclasses of Finsler manifolds – as Minkowski, Berwald, Landsberg, projectively flat and Randers spaces – play a crucial role in such a characterization, see e.g., Bao–Robles–Shen [3], Li [19], and Shen [29, 30, 32].
  • Busemann NPC spaces. Busemann non-positively curved (shortly, NPC) spaces on Riemannian manifolds are characterized by the non-positivity of the sectional curvature, see Busemann [4] and Jost [13]. It was shown however by Kelly–Straus [15] that any strictly convex domain $D\subset\mathbb{R}^{2}$ endowed with the Hilbert metric (which is a projectively flat Finsler manifold with constant negative curvature) is a Busemann NPC space if and only if the boundary of $D$ is an ellipse and the Hilbert geometry is hyperbolic. Ivanov–Lytchak [12] proved that a Finsler manifold is a Busemann NPC space if and only if it is a Berwald manifold of non-positive flag curvature.
  • Sobolev spaces over Finsler manifolds. Contrary to the Riemannian setting, the Sobolev spaces over noncompact, irreversible Finsler manifolds should not have even a vector space structure, see Kristály–Rudas [18]. This construction was carried out over the Funk metric defined on the $n$-dimensional Euclidean unit ball $\mathbb{B}^{n}$.
@@
Let $(M,F,\operatorname{\mathfrak{m}})$ be a Finsler metric measure manifold ($\operatorname{\mathsf{FMMM}}$, for short), i.e., a Finsler manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$, see Section 2 for details. Huang–Kristály–Zhao [10] already pointed out that three non-Riemannian quantities strongly influence the validity of functional inequalities on $(M,F,\operatorname{\mathfrak{m}})$, namely,
@@
  • flag curvature;
  • $S$-curvature, induced by the measure $\operatorname{\mathfrak{m}}$.
@@
It is easy to see that $\lambda_{F}(M)\geq 1$ with equality if and only if $F$ is reversible (i.e., symmetric). Obviously, all Riemannian metrics are reversible, but there are infinitely many irreversible Finsler metrics. Typical examples with infinite reversibility are Funk metric spaces, which served to prove that Sobolev spaces over such Finsler manifolds are not linear structures, see Farkas–Kristály–Varga [9] and Kristály–Rudas [18]. Even if the Finsler metric is explicitly given, it is usually a hard task to determine whether the reversibility is finite or not.
where $R_{y}$ is the Riemannian curvature of $F.$ Clearly, the flag curvature naturally defines the Ricci curvature $\bf Ric$}. As expected, on any Riemannian manifold the flag curvature reduces to the sectional curvature.
Unlike in the Riemannian case, there is no canonical measure on Finsler manifolds. Hence, various measures can be introduced whose behavior may be genuinely different. Every smooth positive measure $\operatorname{\mathfrak{m}}$ on a Finsler manifold $(M,F)$ induces two non-Riemannian quantities, which are the so-called distortion $\tau$ and $S$-curvature $\mathbf{S}$, respectively, see [28]. More precisely, set $\operatorname{d\mathfrak{m}}:=\sigma(x){\operatorname{d}}x^{1}\wedge\cdots\wedge{\operatorname{d}}x^{n}$ in some local coordinate system $(x^{i})$. For any $y\in T_{x}M\backslash\{0\}$, define
@@
The distortion and $S$-curvature are inherently present in Riemannian metric measure manifolds, albeit manifested in different forms. Indeed, if $(M,g,e^{-f}{\operatorname{d}}{\operatorname{vol}_{g}})$ is a weighted Riemannian metric measure manifold, where $f\in C^{\infty}(M)$ and ${\operatorname{d}}{\operatorname{vol}_{g}}$ is the canonical measure, then a straightforward computation shows that for any $(x,y)\in TM$,
$\tau(x,y)=f(x),\quad\mathbf{S}(x,y)=g_{x}(y,\nabla f(x)).$ (1.1)
Clearly, the $S$-curvature does not vanish unless $f$ is a constant. In the Riemannian setting the $S$-curvature cannot preserve a constant sign unless it vanishes identically, whereas this can be achieved by Finslerian metrics; for more details, see the example of Funk metric space below.
Before stating our main results, we recall the main achievements from [10], which represent the starting points of our present study; in fact, they roughly state that if
@@
[…]
@@
  • Caffarelli–Kohn–Nirenberg interpolation inequality; see [10, Theorem 1.1]. If $0<q<2<p$ and 2<n<2(pq)/(p2)2<n<2(p-q)/(p-2), then for every $u\in C^{\infty}_{0}(M)$ one has
@@
On the other hand, if $(M,F,\operatorname{\mathfrak{m}})$ is an $n$-dimensional reversible $\operatorname{\mathsf{FMMM}}$, and the “complementary” case to (1.2) holds, i.e.,
@@
The above set of results contains in certain sense the genesis of our first main result and roughly shows that the validity/failure of functional inequalities on Finsler manifolds is decided by the sign of the $S$-curvature.
@@
where $k,h$ are two constants with $h>k\geq 0$. Then the following statements hold::
@@
    $inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\rm d}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}{|u|^{p}}\operatorname{d\mathfrak{m}}}=0;$
@@
    $inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\rm d}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}=0;$
@@
    $inf_{u\in C^{\infty}_{0}({M})\backslash\{0\}}\frac{\left(\displaystyle\int_{{M}}F^{*2}({\rm d}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{{M}}r^{2}u^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{{M}}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}=0;$
@@
    the Caffarelli–Kohn–Nirenberg interpolation inequality fails on $(M,F,\operatorname{\mathfrak{m}})$, i.e., if $0<q<2<p$ and 2<n<2(pq)/(p2)2<n<2(p-q)/(p-2), then
    $inf_{u\in C^{\infty}_{0}({M})\backslash\{0\}}\frac{\left(\displaystyle\int_{{M}}F^{*2}({\rm d}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{{M}}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{{M}}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}=0.$
As intuitively described by Shen [31], the $S$-curvature describes the change of measure along geodesics. As far as we know, only a few results are available in the literature concerning the impact of $S$-curvature on global properties of Finsler manifolds. In the light of the results from [10], Theorem 1.1 provides a concluding answer about the validity of functional inequalities in Finsler manifolds, which is captured by the behavior of the $S$-curvature. In particular, all the statements except (i) in Theorem 1.1 will be invalid if the condition $h>0$ is relaxed to $h\geq 0$, showing in this way the optimality of our results. Indeed, every reversible Minkowski space endowed with the Busemann–Hausdorff measure is a Cartan–Hadamard manifold with $\mathbf{K}=0$ and $\mathbf{S}=0$ (i.e., $h=k=0$); then the (sharp) inequalities from (1.3)-(1.5) all are valid, which contradict the conclusions from Theorem 1.1. We also notice that there are large classes of non-Riemannian reversible Minkowski spaces; see Example 6.1.
The proof of Theorem 1.1 proceeds by employing polar coordinates on $M$ with a suitable disintegration of the measure $\operatorname{\mathfrak{m}}$. Combined with the key assumption (1.6) with $h>k\geq 0$, comparison principles and suitably chosen test functions (via appropriate approximation arguments), this setup yields the results. For the proof of (i)-(iii) the Gaussian bubble $u_{\alpha}=-e^{-\alpha r}$ is used, while for (iv) a Talentian bubble is considered.
The Gaussian bubble has been used in [10] and Kristály [17] to prove (i) and (ii), respectively, in the special case of the Funk metric on the Euclidean unit ball $\mathbb{B}^{n}$. More generally, for every non-empty bounded strongly convex domain $\Omega\subset\mathbb{R}^{n}$ (containing the origin $\mathbf{0}$) with smooth boundary, there exists a Finsler metric $F$, which is a Funk metric, and $(\Omega,F)$ is a Cartan–Hadamard manifold with $\mathbf{K}=-1/4$; see for example [19, 32, 27]. Such a pair $(\Omega,F)$ is called a Funk metric space and serves as a model space for the application of Theorem 1.1. In the specific case $\Omega=\mathbb{B}^{n}$ and $\operatorname{\mathfrak{m}}=\operatorname{\mathfrak{m}}_{BH}$, the Funk metric is given explicitly by
@@
as in Shen [32, 27, 28]. For this metric, one has ${\bf Ric}=-(n-1)/4$ and ${\bf S}_{BH}=(n+1)/2$, and hence $h=\frac{n+1}{2(n-1)}>\frac{1}{2}=k>0$. By contrast, the symmetrized metric $\tilde{F}(x,y)=(F(x,y)+F(x,-y))/2,$ is precisely the Riemannian Hilbert/Klein metric on $\mathbb{B}^{n}$ with ${\bf K}=-1$ (so ${\bf Ric}=-(n-1)$) and ${\bf S}=0$. In this reversible setting, the sharp inequalities (1.3)–(1.5) hold, and Theorem 1.1 does not apply.
@@
Obviously, every Funk metric space endowed with the Busemann–Hausdorff measure verifies the assumptions of Theorem 1.2. The additional condition $\mathbf{K}\leq 0$ is mild. In fact, if the flag curvature is bounded from below by a positive number, then the Bonnet–Myers theorem implies the compactness of the manifold, see [2, Theorem 7.7.1], in which case the reversibility is finite. In addition, the assumptions on both the flag and Ricci curvatures can be removed in the case of projectively flat Finsler manifolds, see Theorem 1.3 below.
Before stating the next result, we recall that another important notion in Finsler geometry is the weighted Ricci curvature $\mathbf{Ric}_{N}$ for $N\in[n,\infty]$, whose lower bound $\mathbf{Ric}_{N}\geq K$ for some $K\in\mathbb{R}$ is equivalent to the validity of the CD(K,N)(K,N) condition in the Finsler setting (see Ohta [21, 22, 23]). Although $\mathbf{Ric}_{N}$ incorporates both $\mathbf{Ric}$ and $\mathbf{S}$, it cannot replace the separate conditions in (1.6) or (1.8). Indeed, any reversible Minkowski space with the Busemann–Hausdorff measure satisfies $\mathbf{Ric}_{N}=0$; however, on such a space, neither Theorem 1.1 nor Theorem 1.2 holds (see Subsection 6.2). Furthermore, assuming $\mathbf{Ric}_{N}\geq K>0$ is also unsuitable, as this implies the manifold is compact and hence has finite reversibility.
The $S$-curvature also governs the topology of projectively flat Finsler manifolds. A Finsler metric $F$ on a domain $\Omega\subset\mathbb{R}^{n}$ is projectively flat if the images of geodesics are straight lines. The characterization of such metrics on domains in $\mathbb{R}^{n}$ constitutes the regular case of Hilbert’s fourth problem. Although Beltrami’s theorem ensures that every projectively flat Riemannian metric is of constant sectional curvature and vice versa, this is not true for non-Riemannian Finsler metrics, cf. [5, 19, 32]; this fact motivates the study of such Finsler manifolds. The following theorem reveals that a positive lower bound of $S$-curvature influences not only the boundedness and completeness of the domain but also the reversibility of a projectively flat Finsler manifold.
Theorem 1.3.
Let $(\Omega,F,\mathscr{L}^{n})$ be an $n$-dimensional forward complete projectively flat Finsler manifold endowed with the Lebesgue measure. Assume that $\mathbf{S}\geq(n-1)h$ for some constant $h>0$. Then the following statements hold::
  1. (i)
    $\Omega$ is a bounded domain in $\mathbb{R}^{n};$
  2. (ii)
    $(\Omega,F)$ is not backward complete;
  3. (iii)
    $\lambda_{F}(\Omega)=+\infty$.
Note that some statements remain valid under a weaker condition $\mathbf{S}>0$; see subsection 5.2 for details. We also remark that by using $\mathbf{S}\leq-(n-1)h$ instead of $\mathbf{S}\geq(n-1)h$, Theorems 1.11.3 can be established for backward complete Finsler manifolds.
The paper is organized as follows. Section 2 is devoted to preliminaries on Finsler geometry together with some results about volume comparison. In Section 3, we provide some immediate properties of Sobolev spaces over irreversible Finsler manifolds. Here, we also establish an approximation result which will be useful in our main results. The proof of our main Theorem 1.1 is given in Section 4, while the proofs of Theorems 1.2 and 1.3 are presented in Section 5. Section 6 is devoted to the study of Funk metric spaces, by providing examples and counterexamples supporting the sharpness of our main results.